40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Message ID: 1506     Entry time: Sun Sep 7 20:37:50 2014
Author: Tara, Evan 
Type: DailyProgress 
Category: optic 
Subject: Some lens tweaks 

Some minor maintenance/improvements to the optical setup:

  • We replaced the existing lens before the beat PD (RoC = 51.5 mm) with a slightly faster lens (RoC = 38 mm) in order to reduce the spot size on the diode.
  • Tara improved the clamping of mode-matching lenses before the south cavity (they weren't tightened down enough before)
ELOG V3.1.3-